OpenStudy (jadebirdfly):

As punishment for bad behavior during recess, Mrs. Busywork asked her class to multiply 10 by $\frac{1}{3}$ five times. John, however, notices that it is possible to multiply 10 by a single fraction and still get the same answer as the other students. What is this single fraction?

OpenStudy (jadebirdfly):

the fraction is 1/3

Vocaloid (vocaloid):

hint: the teacher wants 10*(1/3)*(1/3)*(1/3)*(1/3)*(1/3) you can multiply all those (1/3)'s together to get a single fraction so what is (1/3)*(1/3)*(1/3)*(1/3)*(1/3)?

Vocaloid (vocaloid):

still there?

OpenStudy (jadebirdfly):

yup thanks i figured it out

OpenStudy (jadebirdfly):

can u help me more?

OpenStudy (jadebirdfly):

@Vocaloid

Vocaloid (vocaloid):

sure, go ahead and post a new question

OpenStudy (jadebirdfly):

If a=3/5 and b=2/3, what is a^2b^-3

OpenStudy (jadebirdfly):

@Vocaloid

Vocaloid (vocaloid):

sorry for the wait if a = 3/5, what is a^2?

OpenStudy (jadebirdfly):

9/25